# Moment bounds for condition numbers and singular values of high-dimensional Gaussian random matrices: Applications and limitations

### v1

@@ 4 Sub-Gaussian Extensions: Limitations and Caveats
Proposition 11 may be viewed as a random matrix example of a phenomenon well known in the general theory of convergence of random variables. While [33, Theorem 5.39] establishes high-probability bounds for the minimum singular value under broad sub-Gaussian assumptions, these bounds govern convergence in probability. Thus, Proposition 11 may be viewed as a random matrix instance where convergence in probability holds but convergence in expectation (and hence moment stability) fails.
We next show that for the sub-Gaussian distribution in Proposition 11, the analogue of Theorem 7 does not hold. In particular, let $X\in\mathbb{R}^{n\times p}$ have i.i.d. entries $\{X_{ij}\}$ supported on $(0,1)$ with distribution function
$F(u)=\frac{1}{\log\!\big(\frac{e}{u}\big)},\qquad u\in(0,1).$ (36)
Define the (uncentered) sample covariance (Gram) matrix
$S:=\frac{1}{n}X^{\top}X,$
and let $\Sigma:=\mathbb{E}[S]=\mathbb{E}[X_{1}X_{1}^{\top}]$, where $X_{1}$ denotes a generic row of $X$. Our next result shows that the expected estimation error of $S^{-1}$, namely $\mathbb{E}\big[\|S^{-1}-\Sigma^{-1}\|_{2}\big]$, also diverges to $\infty$. Consequently, unlike $\mathbb{E}\big[\|S-\Sigma\|_{2}\big]$, it is not possible to obtain a uniform rate for $\mathbb{E}\big[\|S^{-1}-\Sigma^{-1}\|_{2}\big]$ over a general class of sub-Gaussian distributions.
@@
Let $X\in\mathbb{R}^{n\times p}$ have i.i.d. entries with distribution function given in (36), and let $S=\frac{1}{n}X^{T}X$ denote the corresponding sample covariance matrix. Suppose $\lim_{n\rightarrow\infty}\frac{p}{n}=:\gamma$. Then
$\mathbb{E}\big[\|S^{-1}-\Sigma^{-1}\|_{2}\big]\to\infty,$
as $n\to\infty$ for any $\gamma\in[0,1)$.
@@
Observe that
$\|S^{-1}\|_{2}=\frac{1}{\lambda_{\min}(S)}=\frac{1}{\lambda_{\min}\!\left(\frac{1}{n}X^{\top}X\right)}=\frac{n}{s_{\min}(X)^{2}}=\left(\frac{\sqrt{n}}{s_{\min}(X)}\right)^{2}.$
@@
$\mathbb{E}\!\left[\left(\frac{\sqrt{n}}{s_{\min}(X)}\right)^{2}\right]\to\infty,$
as $n\to\infty$ and therefore $\mathbb{E}\|S^{-1}\|_{2}\to\infty$ as $n\to\infty$.
@@
$\mathbb{E}\big[\|S^{-1}-\Sigma^{-1}\|_{2}\big]\geq\mathbb{E}\|S^{-1}\|_{2}-\|\Sigma^{-1}\|_{2}\to\infty,$
$n\to\infty$ which completes the proof.
@@ 4.1 Connection to small-ball probabilities for the smallest singular value
To understand why our results, which hold in the Gaussian case, do not extend to a general class of sub-Gaussian distributions, it is useful to recall a well-known concept in random matrix theory known as small-ball probabilities; see, e.g., [26, 25, 17, 31]. Roughly speaking, the term small-ball probability refers to lower bounds of the form
$\mathbb{P}\big(|\langle X,u\rangle|\geq\epsilon\big)\geq c\qquad\text{for all }u\in\mathcal{S}^{p-1},$
for some fixed constants $\epsilon,c>0$. Intuitively, it quantifies how unlikely it is for a random variable (or a linear functional of a random vector) to fall inside a small neighborhood of zero. Such bounds play a fundamental role in controlling the smallest singular value of random matrices, since excessive mass near zero may allow certain columns or directions to become nearly degenerate and thus force $s_{\min}(X)$ to be extremely small.
To connect our results to the existing literature, we recall [26, Theorem 1.1], stated below in our notation.
Theorem 13 (A small-ball bound for the smallest singular value of sub-Gaussian random matrices)
Let $Z\in\mathbb{R}^{n\times p}$ with $n\geq p$ have i.i.d. entries which are independent copies of a centered sub-Gaussian random variable with unit variance. Then there exist constants $C,c>0$, depending only (polynomially) on the sub-Gaussian norm $\|Z_{11}\|_{\psi_{2}}$, such that for every $\varepsilon>0$,
$\mathbb{P}\!\left(s_{\min}(Z)\leq\varepsilon\big(\sqrt{n}-\sqrt{p-1}\big)\right)\leq(C\varepsilon)^{\,n-p+1}+\exp\{-cn\}.$
Theorem 13 shows that for general i.i.d. sub-Gaussian designs, the lower tail of $s_{\min}(Z)$ admits at best a polynomial decay in $\varepsilon$, up to an additional exponential term $\exp\{-cn\}$ which does not depend on $\varepsilon$. In the Gaussian case, our Theorem 3 (see in particular (2.4)) yields a polynomial-type bound of the same flavor, but importantly does not require the extra exponential term.
This distinction is crucial. Indeed, as emphasized in [26], the presence of the term $\exp\{-cn\}$ is unavoidable for general sub-Gaussian settings. To see this, consider the case where $Z$ has i.i.d. Rademacher entries, taking values $\pm 1$ with equal probability. Then the first two columns of $Z$ coincide with probability
@@
since, for each of the $n$ rows, the two corresponding entries agree with probability $1/2$, and the rows are independent. On this event, the matrix has two identical columns and is therefore rank-deficient, which implies $s_{\min}(Z)=0$. Hence,
$\mathbb{P}\big(s_{\min}(Z)=0\big)\geq 2^{-n+1}=\exp\{-c^{\prime}n\}$
@@
It is also important to note that the exponential term in Theorem 13 is not merely a technical artifact required to accommodate discrete distributions. In fact, in Proposition 11, we construct a continuous sub-Gaussian distribution for which the lower tail of $s_{\min}(Z)$ near the origin decays substantially more slowly, at a logarithmic–polynomial rate - leading to the explosion of moments of $\frac{\sqrt{n}}{s_{\min}(Z)}$ as $n$ increases. This demonstrates that even within the class of continuous sub-Gaussian distributions, it is not guaranteed that the small-ball behavior of $s_{\min}(Z)$ exhibits a purely polynomial decay in $\varepsilon$.
One of our key contributions is to show that under Gaussianity, this additional term can be avoided. This is essential for our main objective of obtaining finite moments for inverse quantities such as $\big(\sqrt{n}/s_{\min}(Z)\big)^{r}$ and $\kappa(Z)^{r}$.
Intuitively, the divergence in Proposition 11 is a consequence of the fact that the entries have an unbounded density near the origin, placing substantial mass close to zero. This increases the probability that certain columns become abnormally small, forcing the smallest (non-zero) singular value to be extremely small and leading to divergence of inverse moments and condition number moments. This leads to another interesting question: can our results, namely Theorem 3 and Theorem 1, be extended to a broad class of sub-Gaussian distributions whose densities are bounded in a neighborhood of the origin? While this appears plausible at an intuitive level, it is not immediate from the existing literature, nor from the structure of our proof of Theorem 3.
To see the source of the difficulty, recall the key step in the proof of Theorem 3 (see (2.4)–(12)). Let $W_{n}=X^{\top}X$, where $X$ has i.i.d. $\mathcal{N}(0,1)$ entries. A crucial Gaussian feature is that for any fixed $u\in\mathcal{S}^{p-1}$,
@@
This identity allows us to convert the problem of bounding $s_{\min}(X)$ into a problem of controlling $\lambda_{\max}(W_{n}^{-1})$, since $\lambda_{\max}(W_{n}^{-1})=\sup_{\|u\|_{2}\leq 1}u^{\top}W_{n}^{-1}u$. The latter is amenable to a standard covering argument: the supremum over the unit ball can be reduced to a supremum over a finite $\varepsilon$-net, and any tail bound for a fixed quadratic form $u^{\top}W_{n}^{-1}u$ can be upgraded to a uniform bound over the net via a union bound. In other words, once one has strong control for each fixed direction $u$, the global supremum can be controlled at essentially the same scale, up to the net cardinality.
For a general sub-Gaussian design, even if the entry distribution has a density bounded near zero, the exact inverse-$\chi^{2}$ distribution of $u^{\top}W_{n}^{-1}u$ is no longer available. As a result, we cannot leverage the same direct route to a sharp control of $\lambda_{\max}(W_{n}^{-1})$. Instead, existing approaches for sub-Gaussian matrices typically proceed by constructing an appropriate good set on which one can control the relevant quadratic forms uniformly (see, e.g., [26, Lemma 7.4 and Lemma 7.6]). The contribution of the complement of this good set is then controlled by a separate argument, and it is precisely at this stage that an additional term of the form $\exp\{-cn\}$ appears. Consequently, even for sub-Gaussian distributions with densities bounded near the origin, it is not clear how one could avoid the extra $\exp\{-cn\}$ term using the above good set approach, and this is exactly the obstruction to obtaining finite inverse moments for the smallest singular value and for the condition number in the same way as in the Gaussian case.
That said, extending our inverse-moment and condition-number results to a natural subclass of sub-Gaussian distributions with bounded densities near the origin remains an interesting and nontrivial problem. We view this as a promising direction for future research building on the present work.

### v2

@@ 4 Sub-Gaussian Extensions: Limitations and Caveats
Proposition 11 may be viewed as a random matrix example of a phenomenon well known in the general theory of convergence of random variables. While [31, Theorem 5.39] establishes high-probability bounds for the minimum singular value under broad sub-Gaussian assumptions, these bounds govern convergence in probability. Thus, Proposition 11 may be viewed as a random matrix instance where convergence in probability holds but convergence in expectation (and hence moment stability) fails.
We next show that for the sub-Gaussian distribution in Proposition 11, the analogue of Theorem 7 does not hold. In particular, let $X_{n}\in\mathbb{R}^{n\times p}$ have i.i.d. entries $\{X_{ij}\}$, where $X_{ij}=Z_{ij}U_{ij}$ for all $(i,j)$, where $\{U_{ij}\}$ and $\{Z_{ij}\}$ are distributed as in the statement of Proposition 11. Define the (uncentered) sample covariance (Gram) matrix
$S_{n}:=\frac{1}{n}X_{n}^{\top}X_{n},$
and let $\Sigma:=\mathbb{E}[S_{n}]$. We have already discussed in Section 3.2 that the positive moments of the estimation error of $S_{n}$, namely $\mathbb{E}\big[\|S_{n}-\Sigma\|_{2}^{r}\big]$, admit a nonasymptotic bound in the Gaussian case (see (30)), which can be extended to the sub-Gaussian setting ( [15]).
Our next result shows, however, that the expected estimation error of the inverse sample covariance, namely $\mathbb{E}\big[\|S_{n}^{-1}-\Sigma^{-1}\|_{2}\big]$, may diverge to $\infty$ for a particular class of sub-Gaussian distributions. Consequently, in contrast to $\mathbb{E}\big[\|S_{n}-\Sigma\|_{2}\big]$, it is not possible to establish a uniform rate for $\mathbb{E}\big[\|S_{n}^{-1}-\Sigma^{-1}\|_{2}\big]$ over a general class of sub-Gaussian distributions.
@@
Let $X_{n}\in\mathbb{R}^{n\times p}$ have i.i.d. entries with distribution function given in (36), and let $S_{n}=\frac{1}{n}X_{n}^{\top}X_{n}$ denote the corresponding sample covariance matrix. Then
$\mathbb{E}\big[\|S_{n}^{-1}-\Sigma^{-1}\|_{2}\big]=\infty,$
for every $n\geq 1$ and $p\leq n$.
@@
Observe that for $p\leq n$
$\|S_{n}^{-1}\|_{2}=\frac{1}{\lambda_{\min}(S_{n})}=\frac{1}{\lambda_{\min}\!\left(\frac{1}{n}X_{n}^{\top}X_{n}\right)}=\frac{n}{s_{\min}(X_{n})^{2}}=\left(\frac{\sqrt{n}}{s_{\min}(X_{n})}\right)^{2}.$
@@
$\mathbb{E}\!\left[\left(\frac{\sqrt{n}}{s_{\min}(X_{n})}\right)^{2}\right]=\infty,$
for every $n\geq 1$ and $p\leq n$. and therefore $\mathbb{E}\|S_{n}^{-1}\|_{2}=\infty$ for every $n\geq 1$..
@@
$\mathbb{E}\big[\|S_{n}^{-1}-\Sigma^{-1}\|_{2}\big]\geq\mathbb{E}\|S_{n}^{-1}\|_{2}-\|\Sigma^{-1}\|_{2}=\infty,$
for every $n\geq 1$ and $p\leq n$.
@@ 4.1 Connection to small-ball probabilities for the smallest singular value
To understand why our results, which hold in the Gaussian case, do not extend to a general class of sub-Gaussian distributions, it is useful to recall a well-known concept in random matrix theory known as small-ball probabilities; see, e.g., [24, 17, 29]. Roughly speaking, the term small-ball probability refers to lower bounds of the form
$\mathbb{P}\big(|\langle X_{n},u\rangle|\geq\epsilon\big)\geq c\qquad\text{for all }u\in\mathcal{S}^{p-1},$
for some fixed constants $\epsilon,c>0$. Intuitively, it quantifies how unlikely it is for a random variable (or a linear functional of a random vector) to fall inside a small neighborhood of zero. Such bounds play a fundamental role in controlling the smallest singular value of random matrices, since excessive mass near zero may allow certain columns or directions to become nearly degenerate and thus force $s_{\min}(X_{n})$ to be extremely small.
To connect our results to the existing literature, we recall [25, Theorem 1.1], stated below in our notation.
Theorem 13 ([25], Theorem 1.1)
Let $Z_{n}\in\mathbb{R}^{n\times p}$ with $n\geq p$ have i.i.d. entries which are independent copies of a centered sub-Gaussian random variable with unit variance. Then there exist constants $C,c>0$, depending only (polynomially) on the sub-Gaussian norm $\|Z_{11}\|_{\psi_{2}}$, such that for every $\varepsilon>0$,
$\mathbb{P}\!\left(s_{\min}(Z_{n})\leq\varepsilon\big(\sqrt{n}-\sqrt{p-1}\big)\right)\leq(C\varepsilon)^{\,n-p+1}+\exp\{-cn\}.$
The above result shows that for general i.i.d. sub-Gaussian designs, the lower tail of $s_{\min}(Z_{n})$ admits at best a polynomial decay in $\varepsilon$, up to an additional exponential term $\exp\{-cn\}$ which does not depend on $\varepsilon$. In the Gaussian case, our Theorem 3 (see in particular (2.4)) yields a polynomial-type bound of the same flavor, but importantly does not require the extra exponential term.
This distinction is crucial. Indeed, as emphasized in [25], the presence of the term $\exp\{-cn\}$ is unavoidable for general sub-Gaussian settings. To see this, consider the case where $Z_{n}$ has i.i.d. Rademacher entries, taking values $\pm 1$ with equal probability. Then the first two columns of $Z_{n}$ coincide with probability
@@
since, for each of the $n$ rows, the two corresponding entries agree with probability $1/2$, and the rows are independent. On this event, the matrix has two identical columns and is therefore rank-deficient, which implies $s_{\min}(Z_{n})=0$. Hence,
$\mathbb{P}\big(s_{\min}(Z_{n})=0\big)\geq 2^{-n+1}=\exp\{-c^{\prime}n\}$
@@
It is also important to note that the exponential term in Theorem 13 is not merely a technical artifact required to accommodate discrete distributions. In fact, in Proposition 11, we construct a continuous sub-Gaussian distribution for which the lower tail of $s_{\min}(Z_{n})$ near the origin decays substantially more slowly, at a logarithmic–polynomial rate - leading to the explosion of moments of $\frac{\sqrt{n}}{s_{\min}(Z_{n})}$ as $n$ increases. This demonstrates that even within the class of continuous sub-Gaussian distributions, it is not guaranteed that the small-ball behavior of $s_{\min}(Z_{n})$ exhibits a purely polynomial decay in $\varepsilon$.
One of our key contributions is to show that under Gaussianity, this additional term can be avoided. This is essential for our main objective of obtaining finite moments for inverse quantities such as $\big(\sqrt{n}/s_{\min}(Z_{n})\big)^{r}$ and $\kappa(Z_{n})^{r}$.
Intuitively, the divergence in Proposition 11 is a consequence of the fact that the entries have an unbounded density near the origin, placing substantial mass close to zero. This increases the probability that certain columns become abnormally small, forcing the smallest singular value to be extremely small and leading to divergence of inverse moments and condition number moments. This leads to another interesting question: can our results, namely Theorem 3 and Theorem 1, be extended to a broad class of sub-Gaussian distributions whose densities are bounded in a neighborhood of the origin? While this appears plausible at an intuitive level, it is not immediate from the existing literature, nor from the structure of our proof of Theorem 3.
To see the source of the difficulty, recall the following key step in the proof of Theorem 3 (see (2.4)–(12)). Let $W_{n}=X_{n}^{\top}X_{n}$, where $X_{n}$ has i.i.d. $\mathcal{N}(0,1)$ entries. A crucial Gaussian feature is that for any fixed $u\in\mathcal{S}^{p-1}$,
@@
This identity allows us to convert the problem of bounding $s_{\min}(X_{n})$ into a problem of controlling $\lambda_{\max}(W_{n}^{-1})$, since $\lambda_{\max}(W_{n}^{-1})=\sup_{\|u\|_{2}\leq 1}u^{\top}W_{n}^{-1}u$. The latter is amenable to a standard covering argument: the supremum over the unit ball can be reduced to a supremum over a finite $\varepsilon$-net, and any tail bound for a fixed quadratic form $u^{\top}W_{n}^{-1}u$ can be upgraded to a uniform bound over the net via a union bound. In other words, once one has strong control for each fixed direction $u$, the global supremum can be controlled at essentially the same scale, up to the net cardinality.
For a general sub-Gaussian design, even if the common distribution of the entries has a density bounded near zero, the exact inverse-$\chi^{2}$ distribution of $u^{\top}W_{n}^{-1}u$ is no longer available. As a result, we cannot leverage the same direct route to a sharp control of $\lambda_{\max}(W_{n}^{-1})$. Instead, existing approaches for sub-Gaussian matrices typically proceed by constructing an appropriate good set on which one can control the relevant quadratic forms uniformly (see, e.g., [25, Lemma 7.4 and Lemma 7.6]). The contribution of the complement of this good set is then controlled by a separate argument, and it is precisely at this stage that an additional term of the form $\exp\{-cn\}$ appears. Consequently, even for sub-Gaussian distributions with densities bounded near the origin, it is not clear how one could avoid the extra $\exp\{-cn\}$ term using the above good set approach, and this is exactly the obstruction to obtaining finite inverse moments for the smallest singular value and for the condition number in the same way as in the Gaussian case.
That said, extending our inverse-moment and condition-number results (Theorem 3 and Theorem 1) to a natural subclass of sub-Gaussian distributions with bounded densities near the origin remains an interesting and nontrivial problem. We view this as a promising direction for future research building on the present work.